OpenStudy (anonymous):

Please help me. In rational expressions, how do we express the following: 1. 729p^3-1 2. 1- 2/x / 1+ 2/x Thanks for the help.

pooja195 (pooja195):

Hey there! Welcome to openstudy! consider reading the CoC: http://openstudy.com/code-of-conduct Heres a link that will help you get to know the site better: http://openstudy.com/study#/updates/548c04e1e4b0f2fbdc23dd6c

pooja195 (pooja195):

do u know where to star?

pooja195 (pooja195):

*start

OpenStudy (anonymous):

I am not sure. My teacher did not give a clear formula. @pooja195

pooja195 (pooja195):

im not sure about this one im sorry @TuringTest ?

OpenStudy (turingtest):

in rational expressions you need to write both the numerator and the denominator what is the denominator of 729p^3-1 ?

OpenStudy (anonymous):

It has no denominator. @TuringTest

OpenStudy (turingtest):

all rational expressions have a denominator, but if that denominator is 1 you don't have to write it, since multiplying and dividing by 1 do not change the original expression

OpenStudy (anonymous):

I tried to work on it but I really don't know how to start. Can you give me some guidelines on how to solve it? @TuringTest .

OpenStudy (turingtest):

a "rational expression" is an expression with the denominator written out notice that \(\frac a1=a\), which means that \(a=\frac a1\) writing \(a\) as a rational expression is writing it instead as \(\frac a1\)

OpenStudy (turingtest):

write your expression with the denominator specified from what I said above you should be able to guess what the denominator of 729p^3-1 is

OpenStudy (anonymous):

Okay, thanks @TuringTest I will try to answer this. Thank you so much!

OpenStudy (turingtest):

you are welcome!